This METAFONT files will generate a font which is used in Austrian schools for hand writing. There are two sizes, 10pt and 20pt.

The Austrian School Writing Font differs from german hand writing fonts in many details; e.g. you have to continue lowercase letters at the middle bar after capital letters F,H, and Z.

The font contains all letters including a sharp s, digits, a few typographic characters .,;:()!? german quotes, and dashes. There is just one accent (umlaut). This should be sufficient for german and english texts.

Additionally are some letter variants for ligatures. In normal circumstances you will NOT need any explicit use of them. The ligtables in the font will do the job for you.

The font is based on the newer T1 encoding and can handle ligatures with umlaut accent. All accented letters (also lowercase \"e and \"i) are contained as glyphs.

Umlaut accent can be applied in 3 different ways:

.) (sure one) as e.g. \"a

.) e.g. "a (without \) works without the package german in the text environment, because the ligatures "a,"o,"u,"A,"O,"U,"s,"`,"``,"', and "'' are part of the font. If you have umlaut accents in normal text use the package german and accents are correct in both environments. Just quotes will not work, since the package german assumes OT1-fonts(!). Use \gsqon/\gsqoff or \gqon/\gqoff respectively.

.) Windows texts with ,,,,, and will be treated correct in the text environment, just (sharp-s) will not work. This is not possible for normal (Computer Modern) text (confer these lines).

Most special characters can be used by an extra command:

\i \j  ıȷ  dotless i or j

    \ii \jj

dotless i or j (variant in combination with FHZ)

    \ae \oe \ue

æœ variant umlaut accent lowercase letters in combination with FHZ

    \B \D \I \N \O \Oe
    \P \S \T \V \W

Ø¶§

capital letters with starting small dash

    \r \x

˚ special characters used at the beginning of a word (confer )

\s  final s (normal s has an extra dash )

    \Fe \He \Ze \Pf  ligatures

Files:

oesch.sty  Package oesch
oe.def  Encoding oe
oesch.mf  Metafont file option 10pt
oeschb.mf  Metafont file option 20pt
oesch_m.mf  Metafont main file
lig.mf  ligtables
readme.txt  short introduction (text)
liesmich.txt short introduction (text) in german
beispiel.tex short introduction (TEX), this file
example.tex short introduction (TEX) in german

Installation: Put oesch.sty and oe.def into a directory, where L^{A}TEX will look for them e.g. TeX/LaTex/misc.

Put *.mf into a directory, where MetaFont will look for them e.g. fonts/source/public/misc. Usage: For short examples use\textoesch{Text}. Long texts use \oeschfamily or

\oeschbfamily. Furthermore I recommend to use it with \renewcommand{\baselinestretch}{1.4}.